OpenStudy (anonymous):

Choose the correct sum of the polynomials (3x3 − 5x − 8) + (5x3 + 7x + 3). 8x^3 + 2x + 5 2x^3 − 12x − 11 2x^3 + 12x − 5 8x^3 + 2x − 5

OpenStudy (anonymous):

8x^3+2x-5

OpenStudy (zehanz):

You can add like terms, so any terms with x³ can be added, as can the terms with x.

OpenStudy (zehanz):

@phani09: please do not give a direct answer...

OpenStudy (anonymous):

@ZeHanz is he right ? if he is can u explain ?

OpenStudy (anonymous):

hey u just asked to choose the correct answer....and thats why i did it...

OpenStudy (anonymous):

all i would want is the answer, and then the explaination to back it up. anyone can jus pick an answer coice

OpenStudy (anonymous):

*choice

OpenStudy (zehanz):

OK, so you go and add the terms with x³: 3x³ + 5x³= ... Now add the terms with x: -5x + 7x = ... Now add the constant terms: -8 + 3 = ... Then put everything on one line

OpenStudy (zehanz):

(I got: 8x³+2x-5)

OpenStudy (anonymous):

thank u !

OpenStudy (zehanz):

Welcome!